# Individual causal effect estimation accounting for latent disease state modification among bipolar participants in mobile health studies

### v1

@@ 2 Department of Statistics, Ohio State University, Columbus, OH, 43210, USA
5 Division of Psychotic Disorders, McLean Hospital, Harvard Medical School, Belmont, MA, 02478, USA
@@ 7 Department of Biostatistics, Columbia University, New York, NY, 10032, USA
Severe mental illness serves as a class to highlight the most serious and long-lasting mental illness conditions, including schizophrenia, severe bipolar disorder, and severe major depression. Individuals living with severe mental illness (SMI) often face difficulties accessing care, poorer health outcomes, and reduced life expectancy [1, 2]. However, the treatment of these individuals is complicated by the heterogeneous nature of SMI between patients and disease states, resulting in varied interventional success [3, 4, 5]. Heterogeneity arises not only from differences in the effect of a given intervention between individuals, but also from differences in the disease states one experiences such as mania, depression, and euthymia. Even when two patients have the same diagnosis, this does not guarantee that they both experience the same disease states or that their disease states manifest in similar ways [6, 7]. This heterogeneity necessitates an individualized analysis approach that can contend with complex causal relationships.
@@
Mobile health (mHealth) studies are a growing body of research that rely on digital technology to record exposures and outcomes in real time [8, 9]. This study design provides rich longitudinal data for each participant and can fill in informational gaps between in-person follow-up with patients. It is particularly useful to study severe mental illness (SMI), such as schizophrenia and bipolar disorder, as the mobile health approach allows researchers to capture fluctuations between disease states and details of at home behavior that may not be evident from clinical visits [10]. Specifically, by employing mobile devices, we are able to gather the intensive longitudinal data needed to study potential at-home interventions at the individual level. Here, we consider the Bipolar Longitudinal Study, a cohort study from McLean Hospital which collects mobile health data from 74 patients with Bipolar disorder, schizophrenia, or schizoaffective disorder for up to five years [11, 12].
@@
Our objective with mobile health data is to study causal effects between a given intervention and outcome, where the cause and effect fluctuate and change over time. Due to the heterogeneous nature of individuals living with SMI and the data collected, we employ an N-of-1 approach where participants and their relative causal effects are studied at the individual level, rather than the aggregate [13, 11, 14]. However, we face several challenges in estimating the individual causal effect, including not directly observing the disease state and missing data. Disease state, representing a subject’s current depressive, manic, or euthymic state, likely acts as an effect modifier of the causal effect of at-home interventions on daily health outcomes. Bipolar disorder is characterized by a patient exhibiting mood states of mania, hypomania, and depression [15, 16]. It has been shown that a disease state framework is also helpful to understand the mood and impairment faced by individuals with schizophrenia across time [5, 17]. In both populations, it is understood that disease state can generate heterogeneous intervention effects within an individual [5, 15, 18], thus acting as a modifier. Accurately estimating the causal effect of a given intervention on health outcomes across disease states for individuals living with SMI will inform clinicians when giving recommendations to improve quality of life and disease prognosis. For example, when studying the effect of physical activity on sleep, the impact of physical activity on sleep might be different depending on the presence of a depressive or manic state thus modifying the observed effect. Unfortunately, it is difficult to observe an individual’s disease state directly from mHealth data. Here, we assume that we collect sufficient information reflective of the disease state such as energy expenditure, mood, and symptoms, such that the state itself is latent. Under this hypothesis, to account for the modification by disease state, we first establish the latent disease state from mobile health data, and secondly include this state when estimating the individual causal effect of interest.
@@
There has been extensive work on identifying latent modifiers through latent modeling techniques. Specifically, this framework is common in the heterogeneous treatment effect literature, where a latent modifier is identified to explain the differences in observed effects between individuals in the sample [19, 20, 21, 22, 23, 24, 25]. These latent modeling methods have been used to better understand heterogeneous effects in a range of settings including the impact of private tutoring [19, 20] pediatric kidney transplant effectiveness [21], and the impact of Medicaid enrollment on emergency department utilization [23]. However, these methods rely on the identification of latent subgroups among individuals in the sample. In the context of latent disease states modifying the individual causal effect, we instead aim to identify the latent classes explaining heterogeneous effects across time points for a given individual. Notably, when differences in the impact or direction of a given treatment exist and are undetected, we may spuriously conclude a treatment to be beneficial or harmful by aggregating across all time. Identifying such heterogeneity not only provides richer interpretation of results but allows for more accurate recommendations dependent on the individual’s current disease state.
@@
To account for a latent variable representing disease state which acts as an effect modifier for the effect of interest, we employ a forward and backward algorithm to learn the disease state from our observed multivariate time series. The hidden Markov model (HMM) provides a framework to identify a latent state from its response time series through estimation of the initial state probabilities, a transition matrix, and the response distribution conditional on the latent state [26, 27]. It is a natural choice to identify latent disease states from mHealth data given its design for time series data with a latent multinomial covariate. While an HMM relies on the Markov property of the response conditional on the latent state, we develop an approach that relaxes this assumption within an auto-regressive hidden Markov model framework [28, 29]. Our proposed model also draws from literature on missing data in HMMs [30, 31, 32, 33], and allows for missingness in the outcome through marginalization and recursive multiple imputation of lagged regressors. Notably, hidden Markov models have been employed previously in similar contexts, such as Hulme et al. employ continuous-time hidden Markov models to the ecological momentary assessment data of individuals with schizophrenia to predict transition to a latent alert state [34] and Ren & Barnet use mixed effect hidden Markov modeling to study diurnal rhythms among adolescents with mood disorders [35]. However, the majority of literature for HMMs has the goal of latent state prediction or identification, not causal estimation, and typically involves less complex assumptions about variable relations [36, 37, 38]. Here, we expand on the existing hidden Markov model literature to our situation of interest and evaluate the performance when estimating an individual causal effect. Specifically, we develop an estimation strategy leveraging the counterfactual approach for causal inference and the HMM for estimation of the individual causal effect and inference.
@@
In the remainder of this paper, we will first define our causal estimand of interest and its relative g-formula. Next, we present a summary of traditional notation and algorithms for constructing an auto-regressive hidden Markov model and explain our adaptations to allow us to fit the model to mobile health data. To evaluate our proposed methods, we compare them in simulations, in application to the Bipolar Longitudinal Study leveraging both actively and passively recorded data. Lastly, we discuss the strengths and limitations of our work, and the future directions needed.
@@ 1 Causal framework
We assume a follow-up length of $T$, and let $t$ index across days of follow-up, such that $t=1,...,T$. We denote the outcome of interest as $Y_{t}$, exposure as $A_{t}$, observed confounders that are impacted by the latent state, such as reported symptoms, as $C_{t}$, and observed confounders that operate independently of the latent state, such as outside temperature, as $W_{t}$. Lastly, we let $L_{t}$ represent the latent disease state, with $J$ possible values. We maintain a flexible framework, where the outcome and/or exposure may be continuous or discrete. Our initial hypothesized causal structure for two time points is illustrated in a directed acyclic graph (DAG) in Figure 1. This framework could be easily adapted to incorporate additional observed confounders and lagged dependencies. Notably, we impose a strong assumption by not including a directed edge between $A_{t}$ and $C_{t}$, however we omit this line as the directionality of such an edge is unknown and likely bi-directional throughout the day. For example, if $A_{t}$ represents daily texts and $C_{t}$ represents physical activity, it is possible that during a day both variables may affect each other. Thus to avoid imposing an incorrect directional assumption, we omit this edge and include lagged effects between the variables ($A_{t-1}$ to $C_{t}$, and $C_{t-1}$ to $A_{t}$).
@@
Figure 1: Directed Acyclic Graph of hypothesized causal relationships between outcome $Y_{t}$, exposure $A_{t}$, latent modifier $L_{t}$, observed confounders that are impacted by the latent state $C_{t}$, and observed confounders that operate independently of the latent state $W_{t}$. Black lines show causal relations within a time point, while grey lines show effects between time points.
@@ 1.1 Contemporaneous Average Individual Causal Effect
Our main focus is estimating the given effect of an intervention at time $t$ on the contemporaneous outcome measured within the same unit of time. To define said treatment effect of interest, we use potential outcome notation as developed by [39, 40, 41, 14]. We first consider the scenario where we wish to interpret this effect for a given value of the latent disease state. Specifically, we define the contemporaneous conditional average individual causal effect of $A_{t}=a_{1,t}$ versus $A_{t}=a_{0,t}$ on $Y_{t}$ given certain $L_{t}=j$ as
@@
As a secondary question, we may wish to calculate the effect of a past intervention at time $t-k$ on a future outcome at time $t$ for some $k\geq 1$ [45]. To denote this effect, we adopt notation from Boruvka et al. [45]. For $k=1$ this effect will be comprised of direct and indirect effects, while for $k>1$ it will be entirely indirect. Thus, we wish to condition only on historical variables that are not descendants of the exposure, $A_{t-k}$. For brevity we will denote the observed regressors for $Y_{t}$ as $X_{Y,t}=(Y_{t-1},A_{t},A_{t-1},C_{t},W_{t})$. We first consider the case where we condition on the value of the latent state at time $t-k$, the k-lagged conditional average individual causal effect, defined as:
$E(Y_{t}^{a_{1,t-k}}-Y_{t}^{a_{0,t-k}}|X_{Y,t-k},L_{t-k}=j\backslash A_{t-k}).$ (3)
@@ 2 Auto regressive hidden Markov model
Hidden Markov models are composed of two variable types: the observed response data ($Y_{t}$) and the latent categorical $L_{t}$ [26, 27]. In a traditional hidden Markov model, the response variable ($Y_{t}$) is caused only by the latent state ($L_{t}$), and the latent state ($L_{t}$) is caused only by the previous value of the state ($L_{t-1}$). Thus, one typically assumes that $L_{t}\perp\kern-5.0pt\perp L_{t-k}|L_{t-1}$ for any $k\geq 2$, and $Y_{t}\perp\kern-5.0pt\perp Y_{t-k}|L_{t}$ for any $k\geq 1$. Under these assumptions, there are three types of parameters to characterize the process. First, there is the initial latent state probability, $\pi_{i}=P(L_{1}=i)$, for $i=1,...,M$. Second, the relationship between the prior and current hidden state is characterized using a transition probability matrix A, where $a_{ij}$ which provides the probability of the current state $L_{t}=j$ given previous state $L_{t-1}=i$, i.e. $a_{ij}=P(L_{t}=j|L_{t-1}=i)$. In their simplest traditional form, these probabilities would be constant across time, but they may also be defined as functions of observed time-varying covariates. The last element of a hidden Markov model describes the distribution or probability of the response variable given the current state, $b_{j}(y_{t})=P(Y_{t}=y_{t}|L_{t}=j)$.
The auto-regressive hidden Markov model (ARHMM) adapts the HMM to a context where the Markovian assumption that $Y_{t}\perp\kern-5.0pt\perp Y_{t-k}|L_{t}$ for any $k\geq 1$ is violated [28, 29]. It allows for $Y_{t}$ to depend on up to $k$ lagged values of the response, meaning for some integer $k>1$, $Y_{t}\perp\kern-5.0pt\perp Y_{t-k-1}|L_{t},Y_{t-1},...,Y_{t-k}$. Thus in an ARHMM, the response distribution will depend not only on the current latent state but also on lagged values of the response variable, meaning $b_{j}(y_{t})=P(Y_{t}|L_{t}=j,Y_{t-1},...Y_{t-k})$. Under a frequentist approach, the initial state probabilities, transition probability matrix, and response distribution are often identified using the Baum Welch algorithm, a forward and backward expectation maximization approach [46]. One first recursively for each $i=1,...,J$ calculates the forwards and backward probabilities, namely $\alpha_{i}(t)=P(Y_{1}=y_{1},...,Y_{t}=y_{t},L_{t}=i)$ and $\beta_{i}(t)=P(Y_{t+1}=y_{t+1},...,Y_{T}=y_{T}|L_{t}=i)$, respectively. These probabilities are then used to update the parameters, and the process is repeated until convergence is reached. Following convergence, researchers commonly employ the Viterbi algorithm to recursively backwards identify the most plausible sequence of the latent state given the parameter estimates [47].
@@ 3 Model specification
As seen in Figure 1, the hypothesized causal structure includes complex relationships and feedback loops among time-varying observed and unobserved variables. Thus the existing hidden Markov model literature requires numerous adaptations for such a model to be appropriate for estimating the causal relationships of interest. First, the model must include three sets of dependent variables to be modeled as descendants of $L_{t}$: $Y_{t},A_{t},$ and $C_{t}$. Each of these time series will require distributional assumptions and parameters to be estimated within the model. We will denote the relative probability distributions given $L_{t}=j$ as $b_{j,Y_{t}}(y_{t}),b_{j,A_{t}}(a_{t}),b_{j,C_{t}}(a_{t})$ for $Y,A,$ and $C$ respectively. We note that within the hidden Markov model literature, these three variables would be labeled as response variables, however we refrain from this terminology, as in our causal framework only $Y_{t}$ is our response of interest, while $A_{t}$ is the intervention, and $C_{t}$ a covariate.
The HMM elements introduced in Section 3 must be defined as functions dependent on covariates of interest. We assume that the initial probabilities are independent of all observed and unobserved variables. However, from Figure 1 we see that for $t>1$, $L_{t}$ depends on not only $L_{t-1}$ but also $C_{t-1},W_{t},Y_{t-1},$ and $A_{t-1}$. Thus the transition probability matrix $\textbf{A}_{t}$, must be defined uniquely at each time point $t$ such that each element is dependent on the causal parents of $L_{t}$. If $J=2$, we use a logit link, and for $J>2$ we use a multinomial logit link. We let
$\eta_{L_{t},ij}=\lambda_{1,ij}+\lambda_{2,ij}C_{t-1}+\lambda_{3,ij}W_{t-1}+\lambda_{4,ij}Y_{t-1}+\lambda_{5,ij}A_{t-1}.C$ (5)
Note that $\eta_{L_{t},ij}$ and $\lambda_{ij}$ only need to be identified for $j>1$, as for $j=1$ the probability will be implied using the link function. Elements $b_{j,Y_{t}}(y_{t}),b_{j,A_{t}}(a_{t}),b_{j,C_{t}}(a_{t})$ must also be estimated not only dependent on $L_{t}$, but also their relative ancestors. Thus, we have:

### v2

@@ 6 Department of Biostatistics, Columbia University, New York, NY, 10032, USA
Severe mental illness serves as a class to highlight the most serious and long-lasting mental illness conditions, including schizophrenia, severe bipolar disorder, and severe major depression. Individuals living with severe mental illness (SMI) often face difficulties accessing care, and have poorer health outcomes and reduced life expectancy [1, 2]. However, the treatment of these individuals is complicated by the heterogeneous nature of SMI between patients and disease states, resulting in varied interventional success [3, 4, 5]. Heterogeneity arises not only from differences in the effect of a given intervention between individuals, but also from differences in the disease states one experiences such as mania, depression, and euthymia. Even when two patients have the same diagnosis, this does not guarantee that they both experience the same disease states or that their disease states manifest in similar ways [6, 7]. This heterogeneity necessitates an individualized analysis approach that can contend with complex causal relationships.
@@
Mobile health (mHealth) studies are a growing body of research that rely on digital technology to record exposures and outcomes in real time [8, 9]. This study design provides rich longitudinal data for each participant and can fill in informational gaps between in-person follow-up with patients. It is particularly useful to study severe mental illness (SMI), such as schizophrenia and bipolar disorder, as the mobile health approach allows researchers to capture fluctuations between disease states and details of at-home behavior that may not be evident from clinical visits [10]. Specifically, by employing mobile devices, we are able to gather the intensive longitudinal data needed to study potential at-home interventions at the individual level. Here, we consider the Bipolar Longitudinal Study, a cohort study from McLean Hospital which collected mobile health data from 74 patients with bipolar disorder, schizophrenia, or schizoaffective disorder for up to five years [11, 12].
@@
Our objective with mobile health data is to study causal effects between a given intervention and outcome, where the cause and effect change over time. Due to the heterogeneous nature of individuals living with SMI and the data collected, we employ an N-of-1 approach where participants and their relative causal effects are studied at the individual level, rather than the aggregate [13, 11, 14]. However, we face several challenges in estimating the individual causal effect, including not directly observing the disease state and missing data. Disease state, representing a subject’s current depressive, manic, or euthymic state, likely acts as an effect modifier of the causal effect of at-home interventions on daily health outcomes. Bipolar disorder is characterized by a patient exhibiting mood states of mania, hypomania, and depression [15, 16]. It has been shown that a disease state framework is also helpful to understand the mood and impairment faced by individuals with schizophrenia across time [5, 17]. In both populations, it is understood that disease state can generate heterogeneous intervention effects within an individual [5, 15, 18], thus acting as a modifier. Accurately estimating the causal effect of a given intervention on health outcomes across disease states for individuals living with SMI will inform clinicians when giving recommendations to improve quality of life and disease prognosis. For example, when studying the effect of physical activity on sleep, the impact of physical activity on sleep might be different depending on the presence of a depressive or manic state thus modifying the observed effect. Unfortunately, it is difficult to observe an individual’s disease state directly from mHealth data. Here, we assume that we collect sufficient information reflective of the disease state such as energy expenditure, mood, and symptoms, such that the state itself is latent. Under this hypothesis, to account for the modification by disease state, we first establish the latent disease state from mobile health data, and secondly include this state when estimating the individual causal effect of interest.
@@
There has been extensive work on identifying latent modifiers through latent modeling techniques. Specifically, this framework is common in the heterogeneous treatment effect literature, where a latent modifier is identified to explain the differences in observed effects between individuals in the sample [19, 20, 21, 22, 23, 24, 25]. These latent modeling methods have been used to better understand heterogeneous effects in a range of settings including the impact of private tutoring [19, 20], pediatric kidney transplant effectiveness [21], and the impact of Medicaid enrollment on emergency department utilization [23]. However, these methods rely on the identification of latent subgroups among individuals in the sample. In the context of latent disease states modifying the individual causal effect, we instead aim to identify the latent classes explaining heterogeneous effects across time points for a given individual. Notably, when differences in the impact or direction of a given treatment exist and are undetected, we may spuriously conclude a treatment to be beneficial or harmful by aggregating across all time. Identifying such heterogeneity not only provides richer interpretation of results but allows for more accurate recommendations dependent on the individual’s current disease state.
@@
To account for a latent variable representing disease state which acts as an effect modifier for the effect of interest, we employ a forward and backward algorithm to learn the disease state from our observed multivariate time series. The hidden Markov model (HMM) provides a framework to identify a latent state from its response time series through estimation of the initial state probabilities, a transition matrix, and the response distribution conditional on the latent state [26, 27]. It is a natural choice to identify latent disease states from mHealth data given its design for time series data with a latent multinomial covariate. While an HMM relies on the Markov property of the response conditional on the latent state, we develop an approach that relaxes this assumption within an auto-regressive hidden Markov model framework [28, 29]. Our proposed model also draws from literature on missing data in HMMs [30, 31, 32, 33], and allows for missingness in the outcome through marginalization and recursive multiple imputation of lagged regressors. Notably, hidden Markov models have been employed previously in similar contexts. Hulme et al. employ continuous-time hidden Markov models to the ecological momentary assessment data of individuals with schizophrenia to predict transition to a latent alert state [34], and Ren & Barnet use mixed effect hidden Markov modeling to study diurnal rhythms among adolescents with mood disorders [35]. However, the majority of literature for HMMs has the goal of latent state prediction or identification, not causal estimation, and typically involves less complex assumptions about variable relations [36, 37, 38]. Here, we expand on the existing hidden Markov model literature to our situation of interest and evaluate the performance when estimating an individual causal effect. Specifically, we develop an estimation strategy leveraging the counterfactual approach for causal inference and the HMM for estimation of the individual causal effect and inference.
@@
In the remainder of this paper, we will first define our causal estimand of interest and its relative g-formula. Next, summarize traditional notation and algorithms for constructing an auto-regressive hidden Markov model and explain our adaptations to allow us to fit the model to mobile health data. To evaluate our proposed methods, we compare them in simulations, in application to the Bipolar Longitudinal Study leveraging both actively and passively recorded data. Lastly, we discuss the strengths and limitations of our work, and the future directions needed.
@@ 1 Causal framework
We assume a follow-up length of $T$, and let $t$ index across days of follow-up, such that $t=1,...,T$. We denote the outcome of interest as $Y_{t}$, exposure as $A_{t}$, observed confounders that are impacted by the latent state, such as reported symptoms, as $C_{t}$, and observed confounders that operate independently of the latent state, such as outside temperature, as $W_{t}$. Lastly, we let $L_{t}$ represent the latent disease state, with $J$ possible values. We maintain a flexible framework, where the outcome and/or exposure may be continuous or discrete. Our initial hypothesized causal structure for two time points is illustrated in a directed acyclic graph (DAG) in Figure 1. This framework could be easily adapted to incorporate additional observed confounders and lagged dependencies. Notably, we impose a strong assumption by not including a directed edge between $A_{t}$ and $C_{t}$ as the directionality of such an edge is unknown and likely bi-directional throughout the day. For example, if $A_{t}$ represents daily texts and $C_{t}$ represents physical activity, it is possible that during a day both variables may affect each other. Thus to avoid imposing an incorrect directional assumption, we omit this edge and include lagged effects between the variables ($A_{t}$ to $C_{t+1}$, and $C_{t}$ to $A_{t+1}$).
@@
Figure 1: Directed Acyclic Graph of hypothesized causal relationships between outcome $Y_{t}$, exposure $A_{t}$, latent modifier $L_{t}$, observed confounders that are impacted by the latent state $C_{t}$, and observed confounders that operate independently of the latent state $W_{t}$. Black lines show contemporaneous causal relations, while grey lines show lagged causal relations.
@@ 1.1 Contemporaneous Average Individual Causal Effect
Our main focus is estimating the given effect of an intervention at time $t$ on the contemporaneous outcome measured within the same unit of time. To define said treatment effect of interest, we use potential outcome notation as developed by [39, 40, 41, 14]. We first consider the scenario where we wish to interpret this effect for a given value of the latent disease state. Specifically, we define the contemporaneous conditional average individual causal effect of $A_{t}=a_{1,t}$ versus $A_{t}=a_{0,t}$ on $Y_{t}$ given $L_{t}=j$ as
@@
As a secondary question, we wish to calculate the effect of a past intervention at time $t-k$ on a future outcome at time $t$ for some $k\geq 1$ [45]. To denote this effect, we adopt notation from Boruvka et al. [45]. For $k=1$ this effect will be comprised of direct and indirect effects, while for $k>1$ it will be entirely indirect. Thus, we wish to condition only on historical variables that are not descendants of the exposure, $A_{t-k}$. For brevity we will denote the observed regressors for $Y_{t}$ as $X_{Y,t}=(Y_{t-1},A_{t},A_{t-1},C_{t},W_{t})$. We first consider the case where we condition on the value of the latent state at time $t-k$, the k-lagged conditional average individual causal effect, defined as:
$E(Y_{t}^{a_{1,t-k}}-Y_{t}^{a_{0,t-k}}|L_{t-k}=j,X_{Y,t-k}\backslash A_{t-k}).$ (3)
@@ 2 Auto regressive hidden Markov model
Hidden Markov models are composed of two variable types: the observed response data ($Y_{t}$) and the latent categorical state $L_{t}$ [26, 27]. In a traditional hidden Markov model, the response variable ($Y_{t}$) is caused only by the latent state ($L_{t}$), and the latent state ($L_{t}$) is caused only by the previous value of the state ($L_{t-1}$). Thus, one typically assumes that $L_{t}\perp\kern-5.0pt\perp L_{t-k}|L_{t-1}$ for any $k\geq 2$, and $Y_{t}\perp\kern-5.0pt\perp Y_{t-k}|L_{t}$ for any $k\geq 1$. Under these assumptions, there are three types of parameters to characterize the process. First, there is the initial latent state probability, $\pi_{i}=P(L_{1}=i)$, for $i=1,...,M$. Second, the relationship between the previous and current hidden state is characterized using a transition probability matrix A, where $a_{ij}$ is the probability of the current state $L_{t}=j$ given previous state $L_{t-1}=i$, i.e., $a_{ij}=P(L_{t}=j|L_{t-1}=i)$. In their simplest form, these probabilities are constants across time, but they may also be defined as functions of observed time-varying covariates. The last element of a hidden Markov model describes the distribution or probability of the response variable given the current state, $b_{j}(y_{t})=P(Y_{t}=y_{t}|L_{t}=j)$.
The auto-regressive hidden Markov model (ARHMM) adapts the HMM to a context where the Markovian assumption that $Y_{t}\perp\kern-5.0pt\perp Y_{t-k}|L_{t}$ for any $k\geq 1$ is violated [28, 29]. It allows for $Y_{t}$ to depend on up to $k$ lagged values of the response, meaning for integer $k>1$, $Y_{t}\perp\kern-5.0pt\perp Y_{t-k-1}|L_{t},Y_{t-1},...,Y_{t-k}$. Thus in an ARHMM, the response distribution will depend not only on the current latent state but also on lagged values of the response variable, meaning $b_{j}(y_{t})=P(Y_{t}|L_{t}=j,Y_{t-1},...Y_{t-k})$. Under a frequentist approach, the initial state probabilities, transition probability matrix, and response distribution are often identified using the Baum Welch algorithm, a forward and backward expectation maximization approach [46]. One first calculates recursively for each $i=1,...,J$ the forward and backward probabilities, namely $\alpha_{i}(t)=P(Y_{1}=y_{1},...,Y_{t}=y_{t},L_{t}=i)$ and $\beta_{i}(t)=P(Y_{t+1}=y_{t+1},...,Y_{T}=y_{T}|L_{t}=i)$, respectively. These probabilities are then used to update the parameters, and the process is repeated until convergence is reached. Following convergence, researchers commonly employ the Viterbi algorithm to recursively backwards identify the most plausible sequence of latent states given the parameter estimates [47].
@@ 3 Model specification
As seen in Figure 1, the hypothesized causal structure includes complex relationships and feedback loops among time-varying observed and unobserved variables. Thus the existing hidden Markov model literature requires numerous adaptations for such a model to be appropriate for estimating the causal relationships of interest. First, the model must include three sets of dependent variables to be modeled as descendants of $L_{t}$: $Y_{t},A_{t},$ and $C_{t}$. Each of these time series will require distributional assumptions and parameters to be estimated within the model. We will denote the relative probability distributions given $L_{t}=j$ as $b_{j,Y_{t}}(y_{t}),b_{j,A_{t}}(a_{t}),b_{j,C_{t}}(c_{t})$ for $Y,A,$ and $C$, respectively. We note that within the hidden Markov model literature, these three variables would be labeled as response variables, however we refrain from this terminology, as in our causal framework only $Y_{t}$ is our response of interest, while $A_{t}$ is the intervention and $C_{t}$ a covariate.
The HMM elements introduced in Section 3 must be defined as functions dependent on covariates of interest. We assume that the initial probabilities are independent of all observed and unobserved variables. However, from Figure 1 we see that for $t>1$, $L_{t}$ depends on not only $L_{t-1}$ but also $C_{t-1},W_{t},Y_{t-1},$ and $A_{t-1}$. Thus the transition probability matrix $\textbf{A}_{t}$ must be defined uniquely at each time point $t$ such that each element is dependent on the causal parents of $L_{t}$. If $J=2$, we use a logit link, and for $J>2$ we use a multinomial logit link. We let
$\eta_{L_{t},ij}=\lambda_{1,ij}+\lambda_{2,ij}C_{t-1}+\lambda_{3,ij}W_{t-1}+\lambda_{4,ij}Y_{t-1}+\lambda_{5,ij}A_{t-1}.$ (5)
Note that $\eta_{L_{t},ij}$ and $\lambda_{ij}$ only need to be identified for $j>1$, as for $j=1$ the probability will be implied using the link function. Elements $b_{j,Y_{t}}(y_{t}),b_{j,A_{t}}(a_{t}),b_{j,C_{t}}(c_{t})$ must also be estimated not only dependent on $L_{t}$, but also their relative ancestors. Thus, we have: